OpenStudy (anonymous):

find the y and x intercept to the equation 2x-10y=-9

OpenStudy (anonymous):

y= 1/5 x +9/10 Go.

OpenStudy (anonymous):

consider the linear equation

OpenStudy (anonymous):

can you write out how you got the answer please

OpenStudy (anonymous):

x-intercept = -4.5 y-intercept = 0.9

OpenStudy (anonymous):

i put rearranged it into a slope intercept form.

OpenStudy (anonymous):

thank you how about -3x-4y=-1

OpenStudy (anonymous):

-3x-4y=-1 -4y=-1+3x y= -3/4 x + 1/4 y-int= 1/4 x-int = 1/3

OpenStudy (anonymous):

how did you get 1/3 for the x-int